Ind AS-Income Expense Items PE with a Cost of . An item of P 100 and a carrying Amount of 80 is revalued toて150. No equ or tax purposes, cumulativeDepreciaton for Tax Purposes is if the item is sold for more than cost - ivalent adjus tment is made 30 and Tax Rate is 30%. Compute DTL to be recognised e) Cumufative Depreciation of 30 (b) Cumulative Depreciation will be included in Taxable Income (taxed at 30%) and the Sale Proceeds will be taxed at 40%. 0 will be included in Taxable Income. Sale Proceeds above Costis not taxable after deducting an Inflation A djusted Cost of 110
Answer #1

A) cost 100....carrying amount 80....revalued figure...150....as per taxation, the cumulative depreciation is 30 and carrying value is 100-30= 70....generally we will conclude that the difference in the tax base is 150-70= 80 and tax rate is 30%...hence differed tax liability is 24...(80*30%)......but here in the question, it is clearly mentioned that revaluation is not allowed as per income tax act.....and any sale proceeds above cost is not taxable.....

Say, the sale is made at a fair value of 150.....then as per income tax...the tax will only be on 100-70..ie., 30 ....tax is 30*30% =9.....and as per books 150-150...profit is 0...and no tax....hence differed tax liablity will always be 9...ie., on the depeeciadepralready provided as per the income tax act.....forfor proce in excess of cost i.e., 50...no differed tax liability is created as in tax no tax is applicable and it will not become temporary difference.....

Hence differed tax liability will be ₹9.

B) in the second case , where sale proceeds were taxed at 40%...after deducting cost inflated of 110₹......same example we will continue....if the asset is sold at 150.....then, as per books pft is 0..(150-150)...as per tax....pft is 150-110=40.....tax on 40 is 40*40%= 16......further, we have to note a point that...as per income tax act....cost is inflated till 110....hence revaluation is allowed till 40...(110-70).......but as per books...revaluation is 70...(150-80).....hence there is a difference in base of 30....(70-40)....on which differed tax of 30% to be levied ..as this has to be removed only through depreciation.....so, differed tax is 30*30%= 9.....

Hence the total differed tax liability to be recognised is 16+9=25₹....

(Additional differed tax on account of revaluation to be recognised in oci...)....

Hope it helps,

Thank u...